One mole nitrogen oxide reacts with oxygen gas to produce a brown gas. Calculate thermochemical equation if 20g of NO completely reacted


1
Expert's answer
2022-06-07T01:39:38-0400


2NO + O2 = 2NO2 ∆H= -114 kJ


2NO + O2 + 114 kJ = 2NO2

60 g NO — 114 kJ

20 g NO — X

X= 20 * 114 / 60 = 38 kJ

Answer: 38 kJ